\begin{thebibliography}{10}

\bibitem{BGAH00}
{\sc C.~Bachoc, T.~A. Gulliver, and M.~Harada}, {\em Isodual codes over
  {$Z_{2k}$} and isodual lattices}, J. Algebraic Combin., 12 (2000),
  pp.~223--240.

\bibitem{BBGS2014}
{\sc A.~Bassa, P.~Beelen, A.~Garcia, and H.~Stichtenoth}, {\em Galois towers
  over non-prime finite fields}, Acta Arith., 164 (2014), pp.~163--179.

\bibitem{BS2019}
{\sc A.~Bassa and H.~Stichtenoth}, {\em Self-dual codes better than the
  {G}ilbert-{V}arshamov bound}, Des. Codes Cryptogr., 87 (2019), pp.~173--182.

\bibitem{BGS05}
{\sc J.~Bezerra, A.~Garcia, and H.~Stichtenoth}, {\em An explicit tower of
  function fields over cubic finite fields and {Z}ink's lower bound}, J. Reine
  Angew. Math., 589 (2005), pp.~159--199.

\bibitem{CPT2020}
{\sc M.~Chara, R.~Podest\'{a}, and R.~Toledano}, {\em Block transitive codes
  attaining the {T}sfasman-{V}ladut-{Z}ink bound}, Des. Codes Cryptogr., 88
  (2020), pp.~1227--1253.

\bibitem{CPQT2024}
{\sc M.~Chara, R.~Podestá, L.~Quoos, and R.~Toledano}, {\em Lifting iso-dual
  algebraic geometry codes}, Des. Codes Cryptogr.,  (2024).

\bibitem{CS94}
{\sc J.~H. Conway and N.~J.~A. Sloane}, {\em On lattices equivalent to their
  duals}, J. Number Theory, 48 (1994), pp.~373--382.

\bibitem{DM03}
{\sc G.~Dorfer and H.~Maharaj}, {\em Generalized {AG} codes and generalized
  duality}, Finite Fields Appl., 9 (2003), pp.~194--210.

\bibitem{GS1995}
{\sc A.~Garc\'{\i}a and H.~Stichtenoth}, {\em A tower of {A}rtin-{S}chreier
  extensions of function fields attaining the {D}rinfeld-{V}l\u{a}du\c{t}
  bound}, Invent. Math., 121 (1995), pp.~211--222.

\bibitem{GS96}
{\sc A.~Garcia and H.~Stichtenoth}, {\em On the asymptotic behaviour of some
  towers of function fields over finite fields}, J. Number Theory, 61 (1996),
  pp.~248--273.

\bibitem{G1981}
{\sc V.~D. Goppa}, {\em Codes on algebraic curves}, Dokl. Akad. Nauk SSSR, 259
  (1981), pp.~1289--1290.

\bibitem{G1982}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Algebraic-geometric
  codes}, Izv. Akad. Nauk SSSR Ser. Mat., 46 (1982), pp.~762--781, 896.

\bibitem{KL2017}
{\sc H.~J. Kim and Y.~Lee}, {\em Construction of isodual codes over {$GF(q)$}},
  Finite Fields Appl., 45 (2017), pp.~372--385.

\bibitem{PSV1991}
{\sc R.~Pellikaan, B.-Z. Shen, and G.~J.~M. van Wee}, {\em Which linear codes
  are algebraic-geometric?}, IEEE Trans. Inform. Theory, 37 (1991),
  pp.~583--602.

\bibitem{Sch89}
{\sc W.~Scharlau}, {\em Selbstduale {G}oppa-{C}odes}, Math. Nachr., 143 (1989),
  pp.~119--122.

\bibitem{S2021}
{\sc L.~Sok}, {\em New families of self-dual codes}, Des. Codes Cryptogr., 89
  (2021), pp.~823--841.

\bibitem{Stich1988}
{\sc H.~Stichtenoth}, {\em Self-dual {G}oppa codes}, J. Pure Appl. Algebra, 55
  (1988), pp.~199--211.

\bibitem{Stich06}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Transitive and
  self-dual codes attaining the {T}sfasman-{V}l\u{a}du\c{t}-{Z}ink bound}, IEEE
  Trans. Inform. Theory, 52 (2006), pp.~2218--2224.

\bibitem{Stichbook09}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Algebraic function
  fields and codes}, vol.~254 of Graduate Texts in Mathematics,
  Springer-Verlag, Berlin, second~ed., 2009.

\bibitem{TVZ1982}
{\sc M.~A. Tsfasman, S.~G. Vl\u{a}du\c{t}, and T.~Zink}, {\em Modular curves,
  {S}himura curves, and {G}oppa codes, better than {V}arshamov-{G}ilbert
  bound}, Math. Nachr., 109 (1982), pp.~21--28.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath, amsthm, amscd, amssymb, amsfonts, amsxtra, amssymb, latexsym}
\usepackage{enumerate}
\usepackage{verbatim}
\usepackage{array}
\usepackage[utf8]{inputenc}
\usepackage{hyperref}
\hypersetup{colorlinks = true,	allcolors  = blue}

\usepackage[svgnames]{xcolor}
\usepackage[normalem]{ulem}

\def\fq{\mathbb{F}_{q}}
\def\fqs{\mathbb{F}_{q^2}}

\def\cC{\mathcal{C}}

\def\cP{\mathbb{P}}

\def\cL{\mathcal{L}}

\def\N{\mathbb{N}}
\def\Z{\mathbb{Z}}

\newcommand{\con}{\operatorname{Con}}

\newcommand{\sop}{\operatorname{Supp}}

\newcommand{\ff}{\operatorname{\mathbb{F}}}

\def\b{\beta}

\def\ff{\mathbb{F}}

\DeclareMathOperator\diff{Diff}

\usepackage{color}

\newcommand\lu[1]{{\color{DarkGreen} #1}}

\newcommand{\sk}{\smallskip}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\newtheorem{remark}[theorem]{Remark}

\begin{document} \sloppy
	\numberwithin{equation}{section}
	\title[Good iso-dual AG-codes from towers of function fields]{Good iso-dual AG-codes from towers \\ of function fields} 
	\author[M.\@ Chara, R.\@ Podestá, L.\@ Quoos, R.\@ Toledano]{María Chara, Ricardo Podestá, Luciane Quoos, Ricardo Toledano}
	\dedicatory{\today}
	\keywords{Iso-dual codes, AG-codes, algebraic function field, tower of function fields} 
	\thanks{2020 {\it Mathematics Subject Classification.} Primary 11T71, 14G50, 94B05, 94B27.} 
	\thanks{The first, second and fourth authors are partially supported by CONICET, FONCyT, SECyT-UNC, and CAI+D-UNL. The third author was financed in part by CAPES-Finance Code 001,  CNPq-Project 307261/2023-9, and FAPERJ - CNE 204.037/2024}	
	\address{María Chara, Researcher of CONICET at Facultad de Ingeniería Química, Universidad Nacional del Litoral, Santiago del Estero 2829, (3000) Santa Fe, Argentina. 
	\newline {\it E-mail: mchara@santafe-conicet.gov.ar}}
	
	\address{Ricardo Podest\'a, FaMAF -- CIEM (CONICET), Universidad Nacional de C\'ordoba, \newline
	Av.\@ Medina Allende 2144, Ciudad Universitaria, (5000) C\'ordoba, Argentina. 
	\newline {\it E-mail: podesta@famaf.unc.edu.ar}}
	
	\address{Luciane Quoos, Universidade Federal do Rio de Janeiro, Centro de Tecnologia, Cidade Universitária, Av.\@ Athos da Silveira Ramos 149, Ilha do Fundão, CEP~21.941-909, Brazil. {\it E-mail: luciane@im.ufrj.br}}
	
	\address{Ricardo Toledano, Facultad de Ingeniería Química, Universidad Nacional del Litoral, Santiago del Estero 2829, (3000) Santa Fe, Argentina. {\it E-mail: ridatole@gmail.com}}
	\dedicatory{\today}

\begin{abstract}
	 	We present a simple method to establish the existence of asymptotically good sequences of iso-dual AG-codes. A key advantage of our approach, beyond its simplicity, is its flexibility, allowing it to be applied to a wide range of towers of function fields. As a result, we present a novel example of an asymptotically good sequence of iso-dual AG-codes over a finite field with 8 elements.
\end{abstract}

\maketitle

\section{Introduction} \label{sec:intro}
Let $\fq$ be a finite field with $q$ elements. A \textit{linear code} $\cC$ over $\fq$ is an $\fq$-linear subspace of $\fq^n$ for $ n\geq 1$. Associated to a linear code $\cC \subseteq \fq^n$ there are three important parameters: its \textit{length} $n$, its \textit{dimension} $k$ as a vector space over $\fq$ and its \textit{minimum} (Hamming) \textit{distance} $d$. 
To emphasize these parameters, it is usual to speak of a linear code $\cC$ as an $[n, k, d]$-code over $\fq$.
The importance of these parameters is that they tell us how good a linear code is for detection and correction of errors. 
In the 1980's, using tools coming from algebraic geometry, Goppa constructed linear codes using algebraic curves defined over a finite field, see \cite{G1981} and \cite{G1982}. The codes constructed by Goppa are usually called algebraic geometry (AG) codes and have played an important role in Coding theory, mainly because Goppa's ideas were used by Tsfasman, Vlăduţ and Zink  in \cite{TVZ1982} to show that there are sequences of linear codes whose parameters are asymptotically better than the Gilbert–Varshamov bound. This certainly was a groundbreaking result at that time. Another important result is that every linear code can be realized as an algebraic geometry code \cite{PSV1991}. 

Let $\cC$ be a linear $[n, k, d]$-code over $\fq$. The ratios 
	$$R(\cC ) = \tfrac{k(\cC )}{n(\cC)}=\tfrac{k}{n} \qquad \text{ and } \qquad \delta (\cC ) = \tfrac{d(\cC)}{n(\cC)}=\tfrac{d}{n},$$ 
 denote the \textit{information rate}
and the \textit{relative minimum distance} of the code, respectively. The set  
\begin{gather*}
U_q = \{ (\delta, R) \in  [0, 1] \times [0, 1] 
\, : \,  \text{there exists a sequence } 
  \{\cC_i \}_{i\geq 0} \text{ of codes over } \fq \\ \text{ such that }  
n(\cC_i) \rightarrow \infty, \delta(\cC_i) \rightarrow \delta, \text{ and } R(\cC_i ) \rightarrow R 
\text{ as } i \rightarrow \infty \},
\end{gather*}
 plays an important role in the asymptotic theory of codes. 
It is well known (see, for example, Proposition~8.4.2 of \cite{Stichbook09}) that there exists a decreasing and continuous function $\alpha_q : [0, 1] \rightarrow [0, 1]$ such that
	\[ U_q = \{(\delta,R) \in  [0, 1] \times [0, 1]  \, : \, 0\leq \delta\leq 1, 0\leq R\leq \alpha_q(\delta) \}. \]
The exact value of $\alpha_q(\delta)$ for $0<\delta<1-q^{-1}$ is still unknown. 

The Gilbert–Varshamov bound is a lower bound of $\alpha_q(\delta)$ for $0<\delta<1-q^{-1}$ which was  improved by the Tsfasman–Vlăduţ–Zink lower bound (TVZ-bound for short) in a certain interval for $q \geq 49$ when $q$ is a square. More precisely (see Theorem 8.4.7 of \cite{Stichbook09}), the 
TVZ-bound establishes that if $\ell$ is a prime power and $q=\ell^2$ then, for any $0 \leq \delta \leq 1-(\sqrt{q}-1)^{-1}$, we have
\begin{equation} \label{eq: TVZ-bound}
	\alpha_q(\delta) \geq \Big(1-\frac{1}{\ell-1} \Big) - \delta.	
\end{equation}	
An infinite sequence $\{\mathcal{C}_i\}_{i\ge 1}$ of $[n_i,k_i,d_i]$-codes over $\ff_q$ is called \textit{(asymptotically) good} if the information rate $R(\mathcal{C}_i) = k_i/n_i$ and the relative minimum distance $\delta(\mathcal{C}_i) = d_i/n_i$ of $\mathcal{C}_i$ satisfy 
 	\[ \liminf_{i \rightarrow \infty} R(\mathcal{C}_i) > 0 \qquad \text{and} \qquad \liminf_{i \rightarrow \infty} \delta(\mathcal{C}_i)>0. \] 

An infinite sequence $\{\mathcal{C}_i\}_{i \ge 1}$ of $[n_i,k_i,d_i]$-codes over $\ff_q$ is said to {\it attain} the
TVZ-bound if, for  $0< \delta\leq 1-(\ell-1)^{-1}$
, we have
	$$ \liminf_{i \rightarrow \infty} R(\mathcal{C}_i) \geq 1- \frac{1}{\ell-1}-\delta \qquad \text{and} \qquad \liminf_{i \rightarrow \infty} \delta(\mathcal{C}_i) \geq  \delta. $$ 

A problem that is still of great interest in coding theory is to construct sequences of codes achieving the TVZ-bound. Significant progress in this direction has been made through the study of sequences of codes derived from towers of algebraic function fields, see for example \cite{Stich06, BBGS2014} and \cite{CPT2020}. A \textit{tower of function fields} over $\fq$ is an infinite sequence $\mathcal{T} = \{F_i\}_{i \ge 0}$ of algebraic function fields $F_i$ over $\fq$ such that $F_{i}$ is a subfield of $F_{i+1}$  for every $i\geq 0$ satisfying
\begin{enumerate}
	\item[($a$)] each extension $F_{i+1}/F_i$ is finite and separable, \sk 
	
	\item[($b$)] $g(F_i)\rightarrow \infty$ as $i\rightarrow\infty$, where $g(F_i)$ denotes the genus of $F_i$, and \sk 
	
	\item[($c$)] $\fq$ is the full constant field of each $F_i$. 
\end{enumerate}

The dual code $\cC^\perp$ of a linear code $\cC$ is the orthogonal complement of $\cC$ in $\fq^n$ with respect to the standard (Euclidean) inner product of $\fq^n$. A code is said to be self-dual if $\cC = \cC^\perp$.
Self-dual AG-codes began to be systematically investigated in \cite{Stich1988}, with generalizations and new constructions proposed in, for example, \cite{DM03} and \cite{S2021}. The good asymptotic behavior of the class of self-dual codes over $\fqs$ was proved for the first time in \cite{Sch89}. This result was later improved in \cite{Stich06}, where it was shown that the classes of transitive iso-dual codes and self-dual codes over $\fqs$ attain the TVZ bound. These results were achieved by considering the Galois closure of a modification of an asymptotically optimal tower over a finite field $\fqs$. Much later, it was shown in \cite{BS2019} that, over almost every non-prime finite field, there are families of self-dual codes that are not only asymptotically good but also asymptotically better than the Gilbert-Varshamov bound for non-prime finite fields of cardinality at least $64$. In this case, the authors used the Galois closure of an asymptotically good tower of function fields over $\ff_q$ studied in \cite{BBGS2014}.

The main goal of this article is to provide a general and relatively simple method to construct asymptotically good sequences of iso-dual AG-codes, which form a slightly more general family of linear codes than the family of self-dual AG-codes. A linear code is called iso-dual if it is equivalent (see Definition \ref{defequiv}) to its dual. Iso-dual codes can be naturally defined not only over finite fields but also over finite rings. For example, they were used in \cite{BGAH00} to study the construction of iso-dual lattices, an interesting class of lattices introduced by Conway and Sloane in \cite{CS94}. One of the main features of our method, besides its simplicity, is its flexibility, in the sense that it can be used in a wide variety of towers of function fields. As a consequence, we are able to prove in Proposition~\ref{prop: TVZ} the existence of a sequence of iso-dual AG-codes over non-prime fields of quadratic cardinality that attains the Tsfasman–Vlăduţ–Zink lower bound in a relatively simple way. We also prove in Proposition~\ref{prop: goodisodualF8} the existence of good sequences of iso-dual AG-codes over finite fields of cubic cardinality and even characteristic, whose lower bound for the relative minimum distance improves the lower bound for the relative minimum distance of the good sequences of self-dual AG-codes given in \cite{BS2019}. In particular, we prove the existence of an asymptotically good sequence of iso-dual AG-codes over $\mathbb{F}_8$. To the best of our knowledge, no example has been given so far of an asymptotically good sequence of either self-dual or iso-dual AG-codes over $\ff_8$.
Finally, it is worth mentioning that our sequences of good iso-dual AG-codes are given in a more explicit way than the constructions of good sequences of self-dual AG-codes given in \cite{Stich06} and \cite{BS2019}. This is because we use explicit recursive towers of function fields instead of the non explicit Galois closure of recursive towers, as it is done in \cite{Stich06} and \cite{BS2019}.

The organization of this paper is as follows. In Section \ref{sec:preliminaries}, we begin by giving the basic definitions and properties we will use from the theory of function fields. Then, we recall the construction of AG-codes using function fields over a finite field and provide the definition of iso-dual codes. We conclude this section with a brief description of the definitions and concepts needed to understand one of the main ingredients of our work: the so-called recursive towers of function fields over a finite field. In Section \ref{sec:lifting}, we recall the second main ingredient of this work, which is the concept of lifting an AG-code in a finite extension of a function field over a finite field. This section is entirely based on our previous work \cite{CPQT2024}. Section \ref{sec: goodsequences} contains our main results. In Theorem \ref{isontowers}, we provide the conditions needed to ensure the existence of good sequences of iso-dual AG-codes using the concept of lifting AG-codes. As a consequence of Theorem \ref{isontowers}, we prove in Proposition \ref{prop: TVZ} the existence of a good sequence of iso-dual AG-codes over a finite field of quadratic cardinality that attains the well-known Tsfasman-Vladut-Zink bound. We then prove in Proposition \ref{prop: goodisodualF8} the existence of good sequences of iso-dual AG-codes over finite fields of cubic cardinality and even characteristic. We conclude this work with Section \ref{succliftings}, where we show that if an iso-dual AG-code $\mathcal{C}$ defined at the bottom field $F_0$ of a tower $\mathcal{T} = \{F_i\}_{i=0}^\infty$ is lifted to an iso-dual AG-code $\mathcal{C}'$ defined at the $n$-th step $F_n$ of the tower, then by lifting $\mathcal{C}$ step by step in the tower until reaching the $n$-th step $F_n$, we obtain the same iso-dual AG-code $\mathcal{C}'$.

\section{Preliminaries} \label{sec:preliminaries}
Here we recall some basic definitions and facts of function fields, divisors, AG-codes and iso-dual AG-codes following the presentation given in \cite{Stichbook09}.

\subsection*{Algebraic function fields} 
Let $\fq$ be a finite field with $q$ elements and let $F$ be an algebraic function field over $\fq$. All function field extensions $F'/F$ considered in this work are always finite, separable and geometric. The last condition means that $\ff_q$ is the full constant field of both $F'$ and $F$.  

We denote by $\mathbb{P}_F$ the set of places in $F$, by $\nu_{P}$ the discrete valuation of $F$ associated to the place $P\in \cP_F$, and by $\mathrm{Div}(F)$ 
the free abelian group generated by the places in $F$. The elements of $\mathrm{Div}(F)$ are called divisors of $F$ and are denoted by $D=\sum_{P \in \mathbb{P}_F} a_P P$, where $a_P=0$ for almost all $P$. The support $\sop(D)$ of a divisor $D$ is the set of places such that $a_P\neq 0$. 

Two divisors $D_1$ and $D_2$ of $F$ are called disjoint divisors if $\sop(D_1)\cap \sop(D_2)= \varnothing$. For a function $z \in F$ we let 
	$$(z)^F, \quad (z)^F_\infty, \quad \text{and} \quad (z)^F_0$$ 
stand for the principal, the pole and the zero divisors of the function $z$ in $F$, respectively. Two divisors $A,B \in \mathrm{Div}(F)$ are equivalent, denoted $A \sim B$, if they differ by a principal divisor, i.e.\@ $B=A+(z)^F$ for some $z \in F$.

Let us consider now a function field extension $F'/F$. For a place $Q$ of $F'$ lying over a place $P$ of $F$, the symbols $e(Q|P)$, $f(Q|P)$ and $d(Q|P)$ denote the \textit{ramification index}, the \textit{inertia degree} and the \textit{different exponent} of $Q$ over $P$, respectively. 

Let $P$ be a place of $F$. It is said that $P$ \textit{splits completely} in $F'$ if there are exactly $[F':F]$ places of $F'$ lying over $P$. This is equivalent to have that $e(Q|P)=f(Q|P)=1$ for every place $Q$ of $F'$ lying over $P$. The opposite situation occurs when there is only one place $Q$ of $F'$ lying over $P$. In this case, if $e(Q|P)=[F':F]$ (hence $f(Q|P)=1$) it is said that $P$ is \textit{totally ramified} in $F'$ and if $f(Q|P)=[F':F]$ (hence $e(Q|P)=1$) it is said that $P$ is \textit{inert} in $F'$.

The \textit{conorm} of a place $P$ of $F$ is the divisor of $F'$ defined as
	$$ \con_{F' /F} (P) = \sum_{Q \in F', Q|P} e(Q|P) \, Q.$$
For a divisor $D\in \mathrm{Div}(F)$ the conorm map defines a group homomorphism from $\mathrm{Div}(F)$ to $\mathrm{Div}(F')$ 
when extended by linearity; 
that is, 
\begin{equation} \label{eq:conorm}
	\con_{F' /F} \Big( \sum_i n_i P_i \Big) = \sum_i n_i \, \con_{F' /F} (P_i).
\end{equation}

The \textit{different divisor} of the extension $F'/F$ is the divisor of $F'$ defined as 
\begin{equation} \label{eq:diff}
	\diff(F'/F) =\sum_{P \in \mathbb{P}_F} \sum_{Q|P} d(Q|P) \, Q,
\end{equation}
where $d(Q|P)$ denotes the different exponent of $Q$ over $P$.
The different divisor is well defined because $d(Q|P)$ is a non-negative integer which is zero  for almost all places.

Finally, we recall that if $g'$ and $g$ denote the genera of $F'$ and $F$ respectively, then
the Riemann-Hurwitz genus formula asserts that 
\begin{equation} \label{eq: genus fla}
	2g'-2 = [F':F](2g-2) + \deg(\diff(F'/F)).
\end{equation}

\subsection*{AG-codes and iso-dual AG-codes} \label{sec:isodual}
Let us recall now the definition of an AG-code. To construct an AG-code over $\fq$, a function field $F$ over $\fq$ and a couple of disjoint divisors of $F$ are needed. More precisely, for a function field $F$ over $\fq$,  we consider two disjoint divisors $D$ and $G$ of $F$ such that 
	$$D=P_1+\dots+P_n,$$ 
where $P_1,\ldots,P_n$ are pairwise distinct rational (i.e.\@ of degree 1) places of $F$. The AG-code $C_{\mathcal{L}}(D, G)$ over $F$ is the linear code over $\fq$ defined as
\begin{equation} \label{defAGcode}
	C_{\mathcal{L}}(D, G)=\{ (z(P_1),\ldots, z(P_n))\, :\, z\in \mathcal{L}(G)\} \subseteq \fq^n,
\end{equation}
where $\cL(G)=\{z\in F\,:\, (z)\ge -G\}\cup \{0\}$ denotes the Riemann-Roch space associated to the divisor $G$. 

%
%
%

Let $\cC \subseteq \fq^n$ be a linear code and ${\bf x} =(x_1, \ldots, x_n) \in (\fq^\ast)^n$. 
The set
	$$ {\bf x} \cdot \cC :=\{ (x_1 c_1, \ldots, x_n c_n) : (c_1, \ldots, c_n) \in \cC \} $$
is clearly another linear code over $\fq$. We notice that the codes $\cC$ and ${\bf x} \cdot \cC$ have the same length, dimension and minimum distance. 
This gives rise to the following notion of equivalence of linear codes as given in Definition 2.2.13 of \cite{Stichbook09}. 
\begin{definition}\label{defequiv}
	Two linear codes $\cC_1$ and $\cC_2$ over $\fq$ are \textit{equivalent} if $\cC_2={\bf x} \cdot\cC_1$ for some ${\bf x} \in (\fq^\ast)^n$. In this case we write $\cC_1 \sim_{\bf x} \cC_2$, or simply $\cC_1 \sim \cC_2$ when the distinction of the vector $\bf x$ is not necessary.
\end{definition}

The dual code $\cC^\perp$ of a linear  code $\cC$ over $\fq$  is the orthogonal complement of $\cC$ in $\fq^n$ with the standard inner product of $\fq^n$. It was shown in Corollary 2.6 of \cite{Stich1988} that the dual of an AG-code is also an AG-code formed with divisors that can be explicitly constructed.

Recall that a linear code $\cC$ is said to be  {\em self-orthogonal} if $\cC \subset \cC^\perp$ and {\em self-dual} if $\cC=\cC^\perp$. The notion of equivalence of linear codes allows to define the following class of linear codes containing the class of self-dual codes.
\begin{definition}
	Let ${\bf x} \in (\fq^*)^n$. A code $\cC \subset \fq^n$ is called \textit{$\bf x$-iso-dual} if $\cC^\perp=\bf x\cdot \cC$. We will simply speak of \textit{iso-dual} codes when there is no need to specify the vector ${\bf x}$.
\end{definition}

Clearly the class of iso-dual codes contains the class of the self-dual codes, that is, $\bf x$-iso-dual codes are just self-dual codes with  ${\bf x}=(1,1,\ldots,1)$. 
We observe that the definition of self-dual code is a particular case of iso-dual codes where the equivalence of codes is given in terms of the monomial equivalence of codes, that is codes which are isometric to their duals via a permutation of coordinates and multiplying certain coordinates by non-zero constants (see, for instance, \cite{KL2017}). 

\subsection*{Towers of function fields}

As it was said in the introduction, a tower $\mathcal{T}$  of function fields  over the finite field $\fq$, is an infinite sequence of function fields $ \{F_i\}_{i \ge 0}$ such that every field extension $F_i/F_{i-1}$ is finite and separable, and such that $\fq$ is the full field of constants of every $F_i$. Also, it is required that the sequences of genera $g(F_i)$ grows to infinity as $i \rightarrow \infty$. We refer to Chapter 7 of \cite{Stichbook09} for more details on definitions and proofs of the results mentioned in this section.

The suitability of a tower $\mathcal{T}=\{F_i\}_{i \ge 0}$ of function fields over $\fq$ in Coding theory applications is determined by the following two  quantities associated to $\mathcal{T}$. The first quantity is the \textit{genus} $\gamma(\mathcal{T})$ of the tower $\mathcal{T}$ over $F_0$, which is defined as 
\begin{equation} \label{eq: genusT}
	\gamma(\mathcal{T}):=\lim_{i \rightarrow \infty}\frac{g(F_i)}{[F_i:F_0]}.
\end{equation} 
 The second quantity is the 
 \textit{splitting rate} $\nu(\mathcal{T})$ of the tower $\mathcal{T}$ over $F_0$, which is defined as
 \[ \nu(\mathcal{T}):= \lim_{i \rightarrow \infty}\frac{N(F_i)}{[F_i:F_0]},\]
 where $N(F_i)$ denotes the number of rational places (or places of degree one) of $F_i$.

These two quantities determine the asymptotic behavior of a tower of function fields over $\ff_q$ as follows. The \textit{limit} of a tower $\mathcal{T}$ over $\ff_q$ is defined as
	\[ \lambda(\mathcal{T}):= \lim_{i \rightarrow \infty} \frac{N(F_i)}{g(F_i)}. \]
It can be proved that $0<\gamma(\mathcal{T})\leq \infty$ and $0\leq \nu(\mathcal{T})<\infty$. Thus, since we have the equation $$\lambda(\mathcal{T})=\frac{\nu(\mathcal{T})}{\gamma(\mathcal{T})},$$ it turns out that $0\leq \lambda(\mathcal{T})<\infty$ for any tower $\mathcal{T}$ over $\ff_q$ and the good or bad asymptotic behavior of $\mathcal{T}$ is decided according to whether or not its limit is zero. 
More precisely, a tower $\mathcal{T}$ of function fields over  $\ff_q$ is said to be \textit{(asymptotically) good} if $\lambda(\mathcal{T})>0$. Otherwise, if $\lambda(\mathcal{T})=0$, the tower is \textit{(asymptotically) bad}. 

Roughly speaking, a tower $\mathcal{T}=\{F_i\}_{i \ge 0}$ over $\ff_q$ is good if $N(F_i)$ grows faster than  the genus of $F_i$ as $i$ goes to infinity. These are the kind of towers of interest in the above mentioned applications. It is well known (see, for instance, Chapter 7 of \cite{Stichbook09}) that if $\mathcal{T}$ is a tower over $\ff_q$, then
\begin{equation}\label{ihara}
	\lambda(\mathcal{T})\leq\frac{1}{\sqrt{q}-1}.
\end{equation}
A tower $\mathcal{T}$ of function fields  over $\ff_q$ is called {\it optimal} if equality in \eqref{ihara} occurs.
 
It was shown by Tsfasman, Vladut and Zink in \cite{TVZ1982}  that for $p$ prime there are optimal towers $\mathcal{T}$ over $\ff_{p^2}$ and $\ff_{p^4}$ constructed with modular curves and Shimura curves. Using these towers the authors proved the existence of AG-codes attaining the TVZ-bound. 
  	 
 In 1995, Garcia and Stichtenoth  
provided in \cite{GS1995} the first explicit construction  of an optimal tower of function fields over $\ff_{q^2}$ using the idea of towers whose function fields are recursively defined by a single equation (see below). This remarkable result opened the door to the explicit construction of AG-codes attaining the TVZ-bound.

 The genus $\gamma(\mathcal{T})$ and the splitting rate $\nu(\mathcal{T})$ of a tower $\mathcal{T}=\{F_i\}_{i \ge 0}$ of function fields over $\fq$ are usually studied in terms of the following two subsets of places of the rational function field $F_0$: 
 the \textit{splitting locus} of $\mathcal{T}$ over $F_0$ is the subset of $ \mathbb{P}_{F_0}$ defined as
	\[ \mathrm{Split}(\mathcal{T}/F_0)=\{P\in \mathbb{P}_{F_0}: \text{$\deg P=1$ and $P$ splits completely in each $F_i$}\}, \]
 while the \textit{ramification locus} of $\mathcal{T}$ over $F_0$ is the subset of $ \mathbb{P}_{F_0}$ defined as
	\[ \mathrm{Ram}(\mathcal{T}/F_0)=\{P\in \mathbb{P}_{F_0}:\text{$P$ ramifies in  $F_i$ for some $i>0$}\}. \]
In many cases  the finiteness of $\mathrm{Ram}(\mathcal{T}/F_0)$ together with the non-emptyness of $\mathrm{Split}(\mathcal{T}/F_0)$ are the two conditions allowing to prove that the tower $\mathcal{T}$ is good. 
More precisely, the inequality $\nu(\mathcal{T})\geq |\mathrm{Split}(\mathcal{T}/F_0)|$ always holds so that $\nu(\mathcal{T})>0$ when $\mathrm{Split}(\mathcal{T}/F_0)\neq \varnothing$ and, in some cases, $\gamma(\mathcal{T})<\infty$ when $|\mathrm{Ram}(\mathcal{T}/F_0)|<\infty$.

All the examples presented later are given in terms of the so called \textit{recursive towers} of function fields, which are the most useful in the applications. Recursive means that the first function field in the tower is a rational function field, that is  $F_0=\fq(x_0)$ for some transcendental element $x_0$ over $\fq$, and all the others fields in the sequence can be defined recursively for every $i\geq 0$ as $F_{i+1}=F_{i}(x_{i+1})$, where $\{x_i\}_{i=0}^{\infty}$ is a sequence of transcendental elements over $\fq$ satisfying an equation of the form $f(x_{i},x_{i+1})=0$, for some appropriate bivariate polynomial $f \in \fq[x,y]$.

\section{Lifting iso-dual AG-codes on function field extensions} \label{sec:lifting}
Let $F'/F$ be an extension of algebraic function fields defined over $\fq$. Next, we recall a construction given recently in \cite{CPQT2024} where an iso-dual AG-code over $F'$ is defined in terms of an iso-dual AG-code over $F$ (see Theorem~\ref{teolevantado} below). That is, starting with an iso-dual AG-code $C_{\cL}(D,G)$ over $F$, an iso-dual AG-code
$C_\cL(\tilde D,\tilde G)$ over $F'$ is constructed. This AG-code over $F'$ is called the \textit{lifting} of $C_{\cL}(D,G)$ to $F'$, and we denote it by
	$$ \mathfrak{L}_{F'/F}(C_\cL(D,G)).$$ 
We can think of $\mathfrak{L}_{F'/F}$ as an operator from the set of AG-codes over $F$ to the set of AG-codes over $F'$. The conditions on $F'$ and $F$ allowing us to define $\mathfrak{L}_{F'/F}(C_\cL(D,G))$ are given in the next theorem whose proof can be found in \cite{CPQT2024}.

\begin{theorem}[\cite{CPQT2024}] \label{teolevantado}
	Let $F'/F$ be a finite separable extension of degree $m\geq 2$ of function fields over $\fq$ with genera $g_{_{F'}} $ and $g_{_F}$, respectively. Let $n, r \in \N$ with $n$ even
	 and suppose that $\{P_1, \dots, P_n\}$ and $\{Q_1, \dots, Q_r\}$ are two disjoint set of places of $F$ such that:
	\begin{enumerate}[$(a)$]
		\item $P_1, \dots, P_{n}$ are rational places and  $P_i$  splits completely in $F'/F$ for $1\leq i\leq n$, \sk
		
		\item  the extension $F'/F$ is unramified outside the set $\{Q_1, \dots , Q_{r}\}$, and \sk 
		
		\item $d(R|Q_i)$ is even for each $1\leq i\leq r$ and each place $R$ of $F'$ lying over $Q_i$.
	\end{enumerate} 
	Let $(\b_1, \ldots, \b_r) \in \Z^r$ be a non-zero $r$-tuple, and consider the disjoint divisors of $F$
	$$D= \sum_{i=1}^n P_i  \qquad \text{and} \qquad G=\sum_{i=1 }^r \b_i Q_i,$$ 
where $\deg(G) >0$, and the associated AG-code $C_{\cL}(D,G)$ over $F$.
Then, the disjoint divisors
	$$\tilde D = \text{Con}_{F'/F}(D) \qquad \text{and} \qquad \tilde G = \con_{F'/F}(G) + \tfrac 12 \mathrm{Diff}(F'/F)$$
of $F'$ define an AG-code $C_{\cL}(\tilde D, \tilde G)$ over $F'$ of length $nm$. The AG-code $C_{\cL}(\tilde D, \tilde G)$ over $F'$ is called the lifting to $F'$ of $C_\cL(D,G)$ over $F$  
and we write \[\mathfrak{L}_{F'/F}(C_\cL(D,G))=C_{\cL}(\tilde D, \tilde G).\]
Furthermore, if $C_{\cL}(D, G)$ is an iso-dual AG-code over $F$ then 
$\mathfrak{L}_{F'/F}(C_\cL(D,G))$ is an iso-dual AG-code over $F'$.
\end{theorem}

The following estimates of the parameters of the lifting of an iso-dual AG-code are straightforward to obtain, as can be seen in \cite{CPQT2024}.
\begin{corollary}[\cite{CPQT2024}] \label{parameters}
	Under the same conditions of Theorem \ref{teolevantado}, if $nm> 2g_{_{F'}}-2$ 
	then $\mathfrak{L}_{F'/F}(C_\cL(D,G))$ is  an $[nm, \frac 12nm, d]$-code
	over $\fq$ where  
	$$ d \ge \tfrac 12nm -g_{_{F'}}+1. $$ 
\end{corollary}

\section{Good sequences of iso-dual AG-codes from towers} \label{sec: goodsequences}

In this section we show that the lifting of an iso-dual AG-code to each field of a suitable good tower of function fields, defines a good sequence of iso-dual AG-codes over a given finite field. As stated in the Introduction, this situation will allow us to prove the existence of good sequences of iso-dual AG-codes in a simpler way than the construction of good self-dual AG-codes given in \cite{Sch89}, \cite{Stich06} and \cite{BS2019}.

We present now our main result, which is a
method to produce asymptotically good sequences of iso-dual AG-codes over $\ff_q$ from towers of function fields.

\begin{theorem}\label{isontowers}
Let $\mathcal{T}=\{F_i\}_{i \ge 0}$ be a tower of function fields over $\fq$ of genus 
	\[ \gamma(\mathcal{T}) \le n (\tfrac 12 -\delta), \] 
for some $0<\delta< \frac 12$, where $n\geq 2$ is an even integer. Let $r$ be a positive integer and suppose that  $\{P_1,\ldots,P_n\}$ and  $\{Q_1,\ldots,Q_r\}$ are two disjoint set of places of $F_0$ such that
\begin{enumerate}[$(a)$]
	\item $P_1,\ldots, P_n$ are rational and $\{P_1,\ldots,P_n\}\subset \mathrm{Split}(\mathcal{T}/F_0)$,\sk
	
	\item 	$\mathrm{Ram}(\mathcal{T}/F_0)\subset \{Q_1,\ldots,Q_r\}$,
	 and  \sk
	
	\item $d(R|Q_j)$ is even for each place $R\in \mathbb{P}_{F_i}$ lying over $Q_j$ for all $i>0$ and $1\leq j\leq r$.	
\end{enumerate}
Consider the disjoint divisors of $F_0$ 
	\[ D=P_1 + \cdots + P_n \qquad \text{and} \qquad G=\sum_{1\le j \le r} \b_jQ_j, \] 
for some integers $\b_1,\ldots, \b_r$, and suppose that the AG-code $C_\cL(D,G)$ is iso-dual. Then, the sequence of liftings $\{\mathcal{C}_i\}_{i>0}$, where			
	$$ \mathcal{C}_i= \mathfrak{L}_{F_i/F_0}(C_\cL(D,G)), $$
is an asymptotically good sequence of iso-dual AG-codes over $\fq$. More precisely, 
$\mathcal{C}_i$ 
is an 
iso-dual AG-code over $\fq$ of length $n_i=n[F_i:F]$ for each $i>0$ such that 
\begin{equation}\label{isorates}
\liminf_{i\rightarrow\infty} R(\mathcal{C}_i)= \tfrac{1}{2} \qquad \text{and} \qquad \liminf_{i\rightarrow\infty} \delta(\mathcal{C}_i)\geq \delta>0,
\end{equation}
where $k_i$ and $d_i$ are, respectively, the dimension and the minimum distance of $\mathcal{C}_i$ 
for each $i> 0$.  
\end{theorem}

\begin{proof}
By hypothesis, we have that $C_\cL(D,G)$ is an iso-dual AG-code over $\fq$ on length $n$. On the other hand, since $\{P_1,\ldots,P_n\}\subset \mathrm{Split}(\mathcal{T}/F_0)$ and for each place $R\in \mathbb{P}_{F_i}$ lying over any $Q\in \mathrm{Ram}(\mathcal{T}/F_0)$ the different exponent $d(R|Q)$ is even, we have from Theorem~\ref{teolevantado} that the lifting $\mathfrak{L}_{F_i/F_0}(C_\cL(D,G))$ is an iso-dual AG-code over $F_i$ of length $n_i=n[F_i:F]$, for any $i>0$. 

Now, let $g_i$ be the genus of $F_i$ and let $m_i=[F_i:F_0]$ for each $i\geq 0$.  The sequence
	\[ \Big\{ \frac{g_i-1}{m_i} \Big \}_{i\geq 0} \]
is monotonically increasing (see, for example, Lemma 7.2.3 of \cite{Stichbook09}) and
	\[ \lim_{i\rightarrow\infty}\frac{g_i-1}{m_i}=\lim_{i\rightarrow\infty} \frac{g_i}{m_i} = \gamma(\mathcal{T}/F_0) \le n (\tfrac{1}{2}-\delta) < \tfrac 12 n. \]
So, we have that there exists $i_0 >0$ such that
	\[ \frac{g_i-1}{m_i} \le n (\tfrac{1}{2}-\delta) < \tfrac{1}{2} n, \]
for each $i\geq i_0$, and thus $n_i=nm_i>2g_i-2$ for \lu{$i\geq i_0$}. Then, from Corollary \ref{parameters} we have
	\[ k_i = \tfrac 12 n_i \qquad \text{and} \qquad d_i \ge \tfrac 12 (n_i-2g_i+2). \]
Therefore,
	$\liminf_{i\rightarrow\infty}R(\mathcal{C}_i) = \frac{1}{2}$ and $ \liminf_{i\rightarrow\infty}\delta(\mathcal{C}_i) \ge \liminf_{i\rightarrow\infty} \Big(\frac{1}{2}-\frac{g_i-1}{n_i}\Big) \ge \delta>0$,
and we conclude that \eqref{isorates} holds.
\end{proof}

In order to use Theorem \ref{isontowers} in a tower $\mathcal{T}=\{F_i\}_{i \ge 0}$, we need to start with an iso-dual AG-code formed with the divisors $D$ and $G$ of $F_0$. There is a particular simple condition ensuring the existence of such a code when the bottom field $F_0$ of $\mathcal{T}$ is a rational function field. More precisely we have the following result. 

\begin{corollary}\label{isoratexamples}
Suppose that $\mathcal{T}=\{F_i\}_{i \ge 0}$ is a tower of function fields over $\ff_q$ satisfying the conditions  
of Theorem \ref{isontowers}, where $F_0=\ff_q(x)$. Consider the disjoint divisors of $F_0$
	\[ D=P_1 + \cdots + P_n \qquad \text{and} \qquad G=\sum_{1 \le j \le r} \b_jQ_j \] 
and the associated AG-code $C_{\cL}(D,G)$. 
If the integers $\b_1,\ldots,\b_r$ are such that
	\[ \sum_{1\le i \le r} \b_i\deg Q_i = \tfrac 12 (n-2),\]
then the sequence $\{\mathcal{C}_i\}_{i > 0} = \mathfrak{L}_{F_i/F_0}(C_\cL(D,G))$
of liftings defines an asymptotically good sequence of iso-dual AG-codes over $\fq$ with rates satisfying \eqref{isorates}.
\end{corollary}

\begin{proof}
It was shown in \cite{Stich1988} that an AG-code $C_{\cL}(D,G)$ of even length $n \geq 2$, defined as in \eqref{defAGcode} over a rational function field, is iso-dual if and only if $\deg G=\tfrac 12 (n-2)$.
Since we have that $C_\cL(D,G)$  is an iso-dual AG-code over $\fq$, then the conclusion follows at once from Theorem \ref{isontowers}.
\end{proof}

In what follows, we present two explicit constructions of sequences of asymptotically good iso-dual codes from well known  recursive towers of function fields.

\subsection{A sequence of iso-dual AG-codes attaining the TVZ-bound} 
\label{goodisodualq2} 

It is well known that there exists a family of self-dual AG-codes attaining the Tsfasman–Vlăduţ–Zink bound \eqref{eq: TVZ-bound}. This was first proved in \cite{Stich06} by using the Galois closure of an asymptotically optimal tower of function fields over $\ff_{q^2}$ presented in  \cite{GS96}. Since the information rate $R$ in the case of self-dual and iso-dual codes is 
	$$R= \tfrac 12,$$ 
having a sequence $\{\mathcal{C}_i\}_{i\ge 0}$ of $[n_i,k_i,d_i]$-self-dual codes (or iso-dual codes) over $\ff_{q^2}$ attaining the Tsfasman–Vlăduţ–Zink bound means that 
	$$\delta = \tfrac 12 - (q-1)^{-1}$$ 
and so the relative minimum distance $\delta(\mathcal{C}_i)$ of $\mathcal{C}_i$ must satisfy
\begin{equation}\label{mindisisodual}
	\liminf_{i \rightarrow \infty} \delta(\mathcal{C}_i)\geq \delta = \frac 12 - \frac{1}{q-1}.
\end{equation}
   
 We are going to show that using Corollary \ref{isoratexamples} with the same asymptotically optimal 
 tower of function fields over $\mathbb{F}_{q^2}$ given in \cite{GS96} we get, in a simpler way, a sequence of iso-dual AG-codes attaining the Tsfasman–Vlăduţ–Zink bound. 

First, we recall some results about the tower given in \cite{GS96} needed to use Corollary~\ref{isoratexamples}. Let $q$ be prime power and let $\mathcal{T}=\{F_i\}_{i \ge 0}$ be  the tower of function fields over $\mathbb{F}_{q^2}$ recursively defined by  $F_0=\mathbb{F}_{q^2}(x_0)$ and  $F_i=F_{i-1}(x_i)$, where for each $i\geq 1$ we have
\begin{equation}\label{eqtower}
x_i^q+x_i=\frac{x^q_{i-1}}{1+x^{q-1}_{i-1}}.
\end{equation}

For $\alpha \in \fqs$, we denote by $P_\alpha$ the only zero of $x_0-\alpha$ in $F_0$, and by $P_\infty$ the only pole of $x_0$ in $F_0$. Consider now the set 
$$\Omega = \{\alpha\in \mathbb{F}_{q^2}:\alpha^q+\alpha=0\}.$$ 

It was shown in Lemmas 3.3, 3.5 and 3.9 of \cite{GS96} that  
\begin{align*}
	& \textrm{Split}(\mathcal{T}/F_0)=\{P_\alpha\in \mathbb{P}_{F_0}:\alpha\notin \Omega\}, \\[1mm] 
	& \textrm{Ram}(\mathcal{T}/F_0)\subset\{P_\alpha\in  \mathbb{P}_{F_0}:\alpha\in \Omega\}\cup \{P_\infty\}.
\end{align*}	
Also, from Remark 3.8 of \cite{GS96}, the genus of $F_n$ is given by
	$$ g(F_n)=\left\{\begin{array}{ll}
	(q^{\frac{n+1}2}-1)^2, 				 & \quad \text{if $n$ is odd}, \\[1.5mm]
	(q^{\frac{n+2}2}-1)(q^{\frac n2}-1), & \quad \text{if $n$ is even}.
\end{array}\right. $$ 
Since $[F_n:F_0]=q^n$, from \eqref{eq: genusT} we conclude that 
	$$\gamma(\mathcal{T}) \le q.$$
Moreover, the cardinality of the splitting set is 
	$$|\textrm{Split}(\mathcal{T}/F_0)|=q^2-|\Omega|=q^2-q.$$ 
\begin{proposition} \label{prop: TVZ}
There is a sequence $\{\mathcal{C}_i\}_{i \ge 1}$ of $\ff_{q^2}$-ary iso-dual AG-codes attaining the Tsfasman–Vlăduţ–Zink bound. 
\end{proposition}

\begin{proof}
Consider the tower $\mathcal{T}=\{F_i\}_{i \ge 0}$  of function fields over $\mathbb{F}_{q^2}$ recursively defined by the equation \eqref{eqtower}. 

For $n=q^2-q$ and $ \delta=\tfrac{1}{2}-\tfrac{1}{q-1}$, 
we have not only that $n\geq 2$ is even but also that
	\[ \gamma(\mathcal{T}) \le q \le n \left(\tfrac{1}{2}-\delta\right), \] 
for any  $q\geq 4$. Now, for any place $P$ of $F_n$ lying over $P_\alpha$ with $\alpha\in \Omega$ or over $P_\infty$, we have from Lemmas~3.3 and 3.5 of \cite{GS96} that either $P$ is unramified in $F_n/F_{n-1}$ or else $P$ is totally ramified in $F_n/F_{n-1}$ with different exponent $d(P|Q)=2(q-1)$, where $Q=P\cap F_{n-1}$.   By using the well known formula (see Corollary 3.4.12 of \cite{Stichbook09}) 
\[d(P|R)=e(P|Q)d(Q|R)+d(P|Q),\]
where $R=Q\cap F_{n-2}$, we  conclude that both different exponents $d(P|P_\infty)$ and $d(P|P_\alpha)$ are even for any $\alpha\in \Omega$. 

Consider the following divisors in the rational function field $F_0$,
	\[D=\sum_{\alpha\notin \Omega}P_\alpha \qquad \text{and} \qquad 
		G=\sum_{\alpha\in\{\infty\}\cup \Omega}\beta_\alpha P_\alpha,\]
where the integers $\beta_\alpha$ are such that 
	\[ \sum_{\alpha\in\{\infty\}\cup \Omega}\beta_\alpha = \tfrac12 (q^2-q-2). \]
Thus, since $q \geq 4$, we have from Corollary \ref{isoratexamples} that the sequence of liftings 
	$$\mathcal{C}_i = \mathfrak{L}_{F_i/F_0}(C_\cL(D,G))$$ 
is a good sequence of iso-dual AG-codes over $\mathbb{F}_{q^2}$ for $q\geq 4$ with the following rates given by \eqref{isorates}
	\[\liminf_{i\rightarrow\infty} R(\mathcal{C}_i) = \frac{1}{2} \qquad \text{and} \qquad \liminf_{i\rightarrow\infty}\delta(\mathcal{C}_i)\geq \frac{1}{2}-\frac{1}{q-1} .\] 
In particular we see that \eqref{mindisisodual} holds so that the sequence of iso-dual AG-codes $\{\mathfrak{L}_{F_i/F_0}(C_\cL(D,G))\}_{i>0}$ over $\ff_{q^2}$  attains the Tsfasman–Vlăduţ–Zink  bound for any prime power $q\geq 4$.	
\end{proof}

\subsection{An asymptotically good sequence of iso-dual codes in characteristic 2} 

In this section we make use the tower of function fields $\mathcal{T}=\{F_i\}_{i\ge 0}$  given in 2005 by Bezerra, García and Stichtenoth \cite{BGS05} to produce good sequences of iso-dual codes over $\ff_{q^3}$ of even characteristic. We refer to this tower as the BGS-tower.

\begin{proposition} \label{prop: goodisodualF8}
For any prime power $q$, there exists a sequence $\{\mathcal{C}_i\}_{i\ge 0}$ of good iso-dual algebraic geometric $[n_i,k_i, d_i]$-codes over $\ff_{q^3}$ with rates 
\begin{equation}\label{eq: ratesF8}
	\liminf_{i\rightarrow\infty}R(\mathcal{C}_i)= \frac{1}{2} \qquad \text{and} \qquad 
	\liminf_{i\rightarrow\infty}\delta(\mathcal{C}_i) \ge \frac{1}{2}-\frac{q+2}{2q(q^2-1)} .
\end{equation}   
\end{proposition}
\begin{proof}
We first recall the definition of the BGS-tower and some of their properties.	
Let $s\in \N$ and let $q$ be a prime power. Let $\mathcal{T}=\{F_i\}_{i \ge 0}$ be the 
tower over $\ff_{q^s}$ recursively defined by the equation
	\[ \frac{1-y}{y^q}=\frac{x^q+x+1}{x}. \]
That is, $F_0=\ff_{q^s}(x_0)$ and $F_i=F_{i-1}(x_i)$ where, for $i\ge 1$, we have 
	\[ \frac{1-x_i}{x_i^q}=\frac{x_{i-1}^q+x_{i-1}+1}{x_{i-1}}. \] 

It was proved in Proposition 1.3 and Theorem 2.9 of \cite{BGS05}  that $\mathrm{Ram}(\mathcal{T}/F_0)$ is a finite set,  $P_\infty\in \mathrm{Ram}(\mathcal{T}/F_0)$, where $P_\infty$ is the only pole of $x_0$ in $F_0$, and also that the genus of this tower over $\ff_{q^s}$ is
	\[ \gamma(\mathcal{T}) = \frac{q+2}{2(q-1)}. \]
It was also proved in Theorem 3.2 of \cite{BGS05}  that when $s=3$ there are $n=q(q+1)$ rational places $P_{\alpha_1},\ldots,P_{\alpha_n}$ of $F_0$ splitting completely in the tower $\mathcal{T}$. Moreover, it was  shown in Propositions 2.5, 2.6, 2.7 and 2.8 of \cite{BGS05} that the different exponents are always even only when $q$ is even.

Thus, we can apply our method given in Theorem \ref{isontowers} by considering this tower $\mathcal{T}$ over cubic finite fields of even characteristic. For this tower $\mathcal{T}$ over $\ff_{q^3}$ with $q$ 
	even, by taking 
	$$\delta=\frac{1}{2}-\frac{q+2}{2q(q^2-1)}, $$ 
we have not only that $0<\delta< \frac 12$, but also that
	$$\gamma(\mathcal{T}) = \frac{q+2}{2(q-1)}= n(\tfrac{1}{2}-\delta).$$
Therefore by considering the following divisors 
	\[ D=\sum_{1\le i \le n} P_{\alpha_i} \qquad \text{and} \qquad G= \tfrac 12 (n-2) P_\infty \]
of $F_0$, we have 
from Corollary \ref{isoratexamples} that the sequence of liftings  $\{\mathfrak{L}_{F_i/F_0}(C_\cL(D,G))\}_{i>0}$ defines a good sequence of iso-dual AG-codes over $\mathbb{F}_{q^3}$ with rates given by \eqref{isorates}, hence \eqref{eq: ratesF8} holds.
\end{proof}

\begin{remark} 
		The above example is interesting because, for $q$ even, our lower bound for the relative minimum distance given in \eqref{eq: ratesF8} --for the good iso-dual AG-codes over $\ff_{q^3}$ defined in Proposition \ref{prop: goodisodualF8}--  improves the lower bound for the relative minimum distance of the self-dual codes given in Theorem 4.3 of \cite{BS2019}. More precisely, it was proved in Theorem~4.3 of \cite{BS2019} that if $q$ is a prime power and  $r\geq 3$ is odd, then there exists a sequence of self-dual codes over $\ff_{q^r}$ whose relative minimum distance satisfies the inequality
		\begin{equation}\label{infrateBS}
			\liminf_{i\rightarrow\infty}\delta(\mathcal{C}_i)\geq \frac{1}{2} - \gamma(\mathcal{F}),
		\end{equation}
		where $\mathcal{F} = \{F_i\}_{i\ge 0}$ 
		is a Galois tower over $\ff_{q^r}$, studied before in \cite{BBGS2014}, of genus
		\begin{equation}\label{torre2}
			\gamma(\mathcal{F}) \le \frac{1}{2} \left( \frac{1}{q^{\frac{r-1}2}-1} + \frac{1}{q^{\frac{r+1}2}-1} \right).
		\end{equation}
		Thus, when $r=3$ we have from \eqref{infrateBS} and \eqref{torre2} that there exists a sequence of good self-dual AG-codes over $\ff_{q^3}$ whose information rate satisfies the inequality
		\begin{equation}\label{infrateBSr=3}
			\liminf_{i\rightarrow\infty} \delta(\mathcal{C}_i) \geq \frac{1}{2} \left(1-\frac{1}{q-1}+\frac{1}{q^2-1}\right) = \frac{1}{2}-\frac{q}{2(q^2-1)}.
		\end{equation}
		It is easy to check that for $q\geq 2$ the inequality 
		$$\frac{1}{2}-\frac{q+2}{2q(q^2-1)} \geq  \frac{1}{2}-\frac{q}{2(q^2-1)}$$
		holds, so that \eqref{eq: ratesF8} improves \eqref{infrateBS} when $q$ is even, as claimed.
		
		 We also note that for $q=2$ and $r=3$ we have from \eqref{torre2} that 
		$\gamma(\mathcal{F}) \le \frac 23$, so that the lower bound \eqref{infrateBS} for the information rate of the self-dual codes defined in Theorem~4.3 of \cite{BS2019} is meaningless. However we can use Proposition \ref{prop: goodisodualF8} with the tower $\mathcal{T}$ over $\ff_{8}$, that is $q=2$, $n=6$ and $\delta= \frac 16$, to conclude that there exists a good sequence of iso-dual AG-codes over $\ff_8$ with rates 
			\begin{equation}\label{limite}
				\liminf_{i\rightarrow\infty}R(\mathcal{C}_i)= \frac{1}{2} \qquad \text{and} \qquad 
				\liminf_{i\rightarrow\infty}\delta(\mathcal{C}_i) \ge \frac{1}{6}.
		\end{equation}  
	\end{remark}

\section{Final remarks: successive liftings}\label{succliftings}
Suppose that we have a tower $\mathcal{T}=\{F_i\}_{i\ge 0}$ of function fields and an iso-dual AG-code $C_{\cL}(D,G)$ over $F_0$. A possible alternative approach to produce good sequences of iso-dual AG-codes from $\mathcal{T}$ could be by iterating the lifting operator step by step in the tower. 
In this final section we show that, under certain splitting conditions on the places defining $D$, the iso-dual AG-code produced by successive liftings of $C_\cL(D,G)$ to $F_1$ and then to $F_2$ and so on until we have lifted the initial code to $F_n$ is the same as the iso-dual AG-code obtained from lifting $C_{\cL}(D,G)$ from $F_0$ to $F_n$ directly.

For example, in the finite tower 
	$$ F_0\subset F_1\subset F_2 $$ 
of function fields over $\ff_q$ we can proceed as follows: if $C_\cL(D,G)$ is an iso-dual code over $F_0$, under the conditions of Theorem \ref{teolevantado} we have that its lifting $\mathfrak{L}_{F_1/F_0}(C_\cL(D,G))$  is an iso-dual AG-code over $F_1$. If with the places defining the lifted AG-code we are again in the conditions of Theorem \ref{teolevantado} in the extension $F_2/F_1$, then we can lift  $\mathfrak{L}_{F_1/F_0}(C_\cL(D,G))$ to an iso-dual AG-code over $F_2$, that is we can consider 
	\[ \left( \mathfrak{L}_{F_2/F_1}\circ \mathfrak{L}_{F_1/F_0}\right)(C_\cL(D,G))=\mathfrak{L}_{F_2/F_1}\left( \mathfrak{L}_{F_1/F_0}(C_\cL(D,G))\right).\]
For this to happen we clearly need that the rational places $P_1,\ldots,P_n$ of $F_0$ defining the divisor $D$ split completely in $F_2$ and that the extension $F_2/F_1$ is unramified outside the places in the support of 
	$$\tilde G = \con_{F_1/F_0}(G) + \tfrac 12 \mathrm{Diff}(F_1/F_0)$$ 
with even different exponent $d(R|Q)$ for any place $R$ of $F_2$ lying over any place $Q$ in the support of 
$\tilde G$.
This last condition is equivalent to have that the different exponents in the different divisor $\mathrm{Diff}(F_2/F_0)$ are even because
\[d(S|Q)=e(S|R)d(R|Q)+d(S|R),\] 
where $S$ is a place of $F_2$, $R=S\cap F_1$ and $Q=R\cap F_0$. 
In particular, we see that if $C_\cL(D,G)$ is an iso-dual code over $F_0$ and the rational places $P_1,\ldots,P_n$ of $F_0$ defining the divisor $D$ split completely in $F_2$ and the different exponents in the divisors $\mathrm{Diff}(F_1/F_0)$ and $\mathrm{Diff}(F_2/F_1)$ are even, not only we can lift to $F_2$ the iso-dual code $\mathfrak{L}_{F_1/F_0}(C_\cL(D,G))$ over $F_1$ but also we have
\begin{equation}\label{liftingitera}
	\mathfrak{L}_{F_2/F_1}\circ \mathfrak{L}_{F_1/F_0}=\mathfrak{L}_{F_2/F_0}.
\end{equation}
In fact, this is so because $\mathfrak{L}_{F_1/F_0}(C_\cL(D,G))=C_\cL(\tilde{D},\tilde{G})$ where
\[\tilde{D}=\con_{F_1/F_0}(D) \qquad \text{and} \qquad \tilde{G}=\con_{F_1/F_0}(G) + \tfrac{1}{2} \mathrm{Diff}(F_1/F_0),\]
so that 
\begin{align*}
	\mathfrak{L}_{F_2/F_1}\left( \mathfrak{L}_{F_1/F_0}(C_\cL(D,G))\right) =\mathfrak{L}_{F_2/F_1}(C_\cL(\tilde{D},\tilde{G})) =C_\cL \big(\tilde{\tilde{D}},\tilde{\tilde{G}} \big),
\end{align*}
where
\[\tilde{\tilde{D}}=\con_{F_2/F_1}(\tilde{D}) \qquad \text{and} \qquad \tilde{\tilde{G}}=\con_{F_2/F_1}(\tilde{G}) + \tfrac{1}{2}\mathrm{Diff}(F_2/F_1).\]
Now
\[\tilde{\tilde{D}}=\con_{F_2/F_1}(\tilde{D})=\con_{F_2/F_1}(\con_{F_1/F_0}(D))=\con_{F_2/F_0}(D),\]
and by Corollary 3.4.12 of \cite{Stichbook09}, we have that
\begin{align*}
	\tilde{\tilde{G}} &= \con_{F_2/F_1}(\tilde{G}) + \tfrac{1}{2}\mathrm{Diff}(F_2/F_1)\\
		&= \con_{F_2/F_1}\left(\con_{F_1/F_0}(G) + \tfrac{1}{2}\mathrm{Diff}(F_1/F_0)\right) + \tfrac{1}{2}\mathrm{Diff}(F_2/F_1)\\
		&= \con_{F_2/F_0}(G) + \con_{F_2/F_1}\left(\tfrac{1}{2}\mathrm{Diff}(F_1/F_0)\right) + \tfrac{1}{2}\mathrm{Diff}(F_2/F_1)\\
		&= \con_{F_2/F_0}(G) + \tfrac{1}{2} \left(\con_{F_2/F_1}\left(\mathrm{Diff}(F_1/F_0)\right) + \mathrm{Diff}(F_2/F_1)\right)\\
		&= \con_{F_2/F_0}(G) + \tfrac{1}{2}\mathrm{Diff}(F_2/F_0),
\end{align*}
so that 
	\[C_\cL \big(\tilde{\tilde{D}},\tilde{\tilde{G}}\big) = \mathfrak{L}_{F_2/F_0}(C_\cL(D,G)),\]
which proves \eqref{liftingitera}.

In particular, equation \eqref{liftingitera} shows that going with the lifting operator step by step in a tower of function fields produces the same sequence of iso-dual AG-codes constructed in Theorem~\ref{isontowers} by lifting directly an iso-dual AG-code over the bottom field to each function field in the tower.

\bibliographystyle{siam}


\end{document}